OpenStudy (anonymous):

(Im not very good with word problems) Together Steve and Tom sold 109 raffle tckets for their school. Steve sold 13 more than 3 times as many raffle tickets as Tom. How many tickets did each boy sell?

OpenStudy (nikato):

let s= # of tickets steve sold and t=# of tickets Tom sold so s+ t=109

Miracrown (miracrown):

We can assume that Steve sold x tickets and Tom sold y tickets. They sold 109 tickets together. That means x+y=109.

OpenStudy (nikato):

steve sold 13 more than 3 times Tom so s= 13+ 3t

OpenStudy (nikato):

substitute this second equation into the first equation 13+3t +t=109 and now solve for t

OpenStudy (anonymous):

got it thanks!

OpenStudy (nikato):

ok. you're welcome